OpenStudy (anonymous):

A cable has to connect between a point 3 distance units south of the origin i.e. (0, −3) and the parabola y^2 = −x. Find the point on the parabola if the cable length is to be minimised.

OpenStudy (amistre64):

d^2 = (x-x1)^2 + (y-y1)^2 might be useful in this

OpenStudy (anonymous):

What do I do with that?

OpenStudy (amistre64):

d' = 2x'(x-x1) + 2y'(y-y1) 0 = x'(x-x1) + y'(y-y1) let x1, y1 be the point; and let x = -y^2, and y = y be the second point x' = -2y y' = 1 0 = -2y(-y^2-0) + (y+3) 0 = 2y^3+y+3 solve for y

OpenStudy (amistre64):

might do better with a different choice of the issues

OpenStudy (amistre64):

the wolf makes it simple. another route is to simply find when the slope between (0,-3) and (-y^2,y) is perp to the tangent of the parabola

OpenStudy (amistre64):